OpenStudy (anonymous):

find the slope of the line through the pair of points. n(5,0), k(-5, -9) A. 10/9 B. 0 C. 9/10 D. - 9/10

OpenStudy (anonymous):

slope= change in y over change in x or slope= y2-y1/x2-x1 so C 9/10

OpenStudy (vshiroky):

-9-0/-5-5=-9/10

OpenStudy (anonymous):

\[slope= \frac{ -9-0 }{ -5-5 }=\frac{ -9 }{ -10 }=\frac{ 9 }{ 10 }\]